OpenStudy (xxanntdmxx):

will medal and fan A circular piece of metal has a radius of 3.5 meters. The metal sells for $3.25 per square meter. What is the total cost of the circular piece of metal? A. $11.38 B. $38.47 C. $71.44 D. $125.01

OpenStudy (anonymous):

Okay so do you know how to find the area of a circle?

OpenStudy (xxanntdmxx):

sadly, no.

OpenStudy (anonymous):

okay thats okay we all forget and theres always time to learn something new so the area for a circle is A= 3.14(pi/exactly)*Radius(3.5)*Radius(3.5)

OpenStudy (anonymous):

do you understand that or did i confuse you?

OpenStudy (xxanntdmxx):

um, kinda

OpenStudy (xxanntdmxx):

i got D. is that correct?

OpenStudy (anonymous):

ok illtake you through this step by step. 1. Area= 3.14*r*r 2. Area= 3.14*3.5*3.5 3.Area= 3.14*12.15 4. Area= 38.64500 5. So we round, 38.64500 could not be a price we round to 2 numbers after the decimal

OpenStudy (anonymous):

my bad 5. so we round 38..465000

OpenStudy (anonymous):

we round to a logical price someone could pay and we get 38.47

OpenStudy (anonymous):

so we get our answer to buy the piece of circular medal with a radius of 3.5 it would be that it would cost 38.47

OpenStudy (anonymous):

I hope that I helped your final answer is B